# PaSa: An LLM Agent for Comprehensive Academic Paper Search

### v1

@@ 5.4 Ablation study
We perform ablation studies in Table 8 to evaluate the individual contributions of exploring citation networks, RL training, and using the Selector as the reward model. The results indicate that removing the [Expand] action from the Crawler leads to a significant drop in the recall: a decrease of 22.98% on AutoScholarQuery and 32.21% on RealScholarQuery. Furthermore, RL training enhances recall by 6.24% on AutoScholarQuery and 19.96% on RealScholarQuery. The RL training curves are depicted in Figure 3, where the training curves show a steady increase in return with the training steps, eventually converging after 200 steps. Finally, removing the Selector as an auxiliary reward model results in a 3.76% recall drop on AutoScholarQuery and a 9.63% drop on RealScholarQuery.
$\alpha$ Crawler Recall Crawler Actions
0.5 0.7227 175.9
1.0 0.7708 319.8
1.5 0.7931 382.4
2.0 0.8063 785.5
@@ 6 Conclusion
In this paper, we introduce PaSa, a novel paper search agent designed to provide comprehensive and accurate results for complex academic queries. PaSa is implemented within the AGILE, a reinforcement learning framework for LLM agents. To train PaSa, we developed AutoScholarQuery, a dataset of fine-grained academic queries and corresponding papers drawn from top-tier AI conference publications. To evaluate PaSa in real-world scenarios, we also constructed RealScholarQuery, a dataset of actual academic queries paired with annotated papers. Our experimental results demonstrate that PaSa outperforms all baselines, including Google, Google Scholar, and Google with GPT-4o, ChatGPT, GPT-o1, and PaSa-GPT-4o. In particular, PaSa-7B surpasses Google with GPT-4o by 37.78% in recall@20 and 39.90% in recall@50, while also exceeding PaSa-GPT-4o by 30.36% in recall and 4.25% in precision. These findings underscore PaSa significantly improves the efficiency and accuracy of academic search.
Appendix A Implementation Details of the Crawler
A.1 Imitation learning data generation
For expand sessions starting from $S_{q+p}$, we continue by searching for the generated queries using Google. We then sample papers from the search results and obtain the initial state, which includes both the query and a paper. To build the session trajectory, we examine each sub-section of the paper. If the sub-section references at least one paper in the AutoScholarQuery training set corresponding to the query, the sub-section is selected. Otherwise, the sub-section is selected with a 10% probability to enhance trajectory diversity. The selected sections are filled into the template in Table 11, completing the session trajectory. In total, 9,978 expand session trajectories are constructed.
The prompt for search query generation.
@@ A.1 Imitation learning data generation
A.2 Roll-Out in PPO training
During PPO training, each device processes 4 user queries in each step, generating a search session for each user query. Then, 6 expansion sessions are created by randomly sampling 6 papers from the search results. This process is repeated with the expand citation results, yielding 6 additional expand sessions. In total, 16 session trajectories are generated per step.
Appendix B Implementation Details of the Selector
We begin by sampling user queries from the AutoScholarQuery training set. For each user query and one of its corresponding papers in the AutoScholarQuery training set, we prompt GPT-4o to generate a decision token and rationale (see Table 15 for prompt). We reject any data where the decision token is "False", as this contradicts the AutoScholarQuery label. The remaining data are retained as positive <user query, paper> pairs.
@@ Appendix B Implementation Details of the Selector
Appendix C Selector Test Dataset
@@
Appendix D Dataset Examples
Table 12 shows the examples of queries and corresponding papers in RealScholarQuery.
Query: Give me papers about how to rank search results by the use of LLM Query Date: 2024-10-01 Answer Papers: [0] Instruction Distillation Makes Large Language Models Efficient Zero-shot Rankers [1] Beyond Yes and No: Improving Zero-Shot LLM Rankers via Scoring Fine-Grained Relevance Labels [2] Large Language Models are Effective Text Rankers with Pairwise Ranking Prompting [3] A Setwise Approach for Effective and Highly Efficient Zero-shot Ranking with Large Language Models [4] RankVicuna: Zero-Shot Listwise Document Reranking with Open-Source Large Language Models [5] PaRaDe: Passage Ranking using Demonstrations with Large Language Models [6] Is ChatGPT Good at Search? Investigating Large Language Models as Re-Ranking Agents [7] Large Language Models are Zero-Shot Rankers for Recommender Systems [8] TourRank: Utilizing Large Language Models for Documents Ranking with a Tournament-Inspired Strategy [9] ExaRanker: Explanation-Augmented Neural Ranker [10] RankRAG: Unifying Context Ranking with Retrieval-Augmented Generation in LLMs [11] Make Large Language Model a Better Ranker [12] LLM-RankFusion: Mitigating Intrinsic Inconsistency in LLM-based Ranking [13] Improving Zero-shot LLM Re-Ranker with Risk Minimization [14] Zero-Shot Listwise Document Reranking with a Large Language Model [15] Consolidating Ranking and Relevance Predictions of Large Language Models through Post-Processing [16] Re-Ranking Step by Step: Investigating Pre-Filtering for Re-Ranking with Large Language Models [17] Large Language Models for Relevance Judgment in Product Search [18] PromptReps: Prompting Large Language Models to Generate Dense and Sparse Representations for Zero-Shot Document Retrieval [19] Passage-specific Prompt Tuning for Passage Reranking in Question Answering with Large Language Models [20] When Search Engine Services meet Large Language Models: Visions and Challenges [21] RankZephyr: Effective and Robust Zero-Shot Listwise Reranking is a Breeze! [22] Rank-without-GPT: Building GPT-Independent Listwise Rerankers on Open-Source Large Language Models [23] MuGI: Enhancing Information Retrieval through Multi-Text Generation Integration with Large Language Models [24] Discrete Prompt Optimization via Constrained Generation for Zero-shot Re-ranker [25] REAR: A Relevance-Aware Retrieval-Augmented Framework for Open-Domain Question Answering [26] Agent4Ranking: Semantic Robust Ranking via Personalized Query Rewriting Using Multi-agent LLM [27] FIRST: Faster Improved Listwise Reranking with Single Token Decoding [28] Leveraging LLMs for Unsupervised Dense Retriever Ranking [29] Unsupervised Contrast-Consistent Ranking with Language Models [30] Enhancing Legal Document Retrieval: A Multi-Phase Approach with Large Language Models [31] Found in the Middle: Permutation Self-Consistency Improves Listwise Ranking in Large Language Models [32] Fine-Tuning LLaMA for Multi-Stage Text Retrieval [33] Zero-shot Audio Topic Reranking using Large Language Models [34] Uncovering ChatGPT’s Capabilities in Recommender Systems [35] Cognitive Personalized Search Integrating Large Language Models with an Efficient Memory Mechanism [36] Towards More Relevant Product Search Ranking Via Large Language Models: An Empirical Study [37] Pretrained Language Model based Web Search Ranking: From Relevance to Satisfaction [38] Open-source large language models are strong zero-shot query likelihood models for document ranking
Table 12: Examples of queries and corresponding papers in RealScholarQuery.
Appendix E Prompt Templates
E.1 Prompts for Baselines
Table 13 exhibits the search query paraphrasing prompt for the baseline model Google with GPT-4o.
Table 14 exhibits the prompt for the baseline model ChatGPT (search-enabled GPT-4o).
The prompt for search query paraphrase.
@@ E.1 Prompts for Baselines
Table 13: The prompt for search query paraphrase.
The prompt for ChatGPT (search-enabled GPT-4o).
@@
Table 14: The prompt for Chatgpt (search-enabled GPT-4o).
E.2 Prompt for Paper Selection
Table 15 shows the prompt for PaSa selector and gpt-4o to judge whether a paper matches the requirements of the user’s query.
The prompt for paper selection.
@@ E.2 Prompt for Paper Selection
Table 15: The prompt used with pasa selector or GPT-4o to judge the selection of the paper.
Table 16 presents the prompt template used with GPT-4o to automatically generate AutoScholarQuery. For each paper, we extract the Related Work section, input it into GPT-4o, and use the prompt to extract scholarly queries and their corresponding paper answers from the Related Work section.
The prompt for AutoScholarQuery generation.
[…]
Table 16: The prompt used with GPT-4o to automatically generate AutoScholarQuery.

### v2

@@ 5.3 Main results
We observe that using multiple ensembles of Crawler during inference can improve performance. Specifically, we use sampling decoding to run Crawler twice in the PaSa-7b-ensemble setting, which boosts Crawler recall by 3.34% on AutoScholarQuery and increases the final recall by 1.51%, with no significant change in precision.
As both the final decision-maker and auxiliary reward model in RL training for the Crawler, the performance of the Selector is crucial. To evaluate its effectiveness, we collected a dataset of 200 query-paper pairs, annotating whether each paper meets the query’s requirements. This dataset serves as the benchmark for evaluating the Selector (see Appendix F for details). We then compared our Selector against GPT-4o Hurst et al. (2024) and Qwen-2.5-7b Yang et al. (2024), as shown in Table 6. The results show that our Selector achieves an F1 score of 85%, outperforming GPT-4o by 5% and Qwen-2.5-7b by 30%. Additionally, when compared to a setting where reasoning precedes decision token generation, the performance is comparable. Lastly, the Selector’s precision reaches 95%, confirming its effectiveness as an auxiliary reward model for the Crawler RL training.
@@ 5.4 Ablation study
We perform ablation studies in Table 7 to evaluate the individual contributions of exploring citation networks, RL training, and using the Selector as the reward model. The results indicate that removing the [Expand] action from the Crawler leads to a significant drop in the recall: a decrease of 22.98% on AutoScholarQuery and 32.21% on RealScholarQuery. Furthermore, RL training enhances recall by 6.24% on AutoScholarQuery and 19.96% on RealScholarQuery. The RL training curves are depicted in Figure 3 in Appendix D.2, where the training curves show a steady increase in return with the training steps, eventually converging after 200 steps. Finally, removing the Selector as an auxiliary reward model results in a 3.76% recall drop on AutoScholarQuery and a 9.63% drop on RealScholarQuery.
$\alpha$ Crawler Crawler Precision Recall
Recall Actions
0.5 0.7227 175.9 0.1458 0.4602
1.0 0.7708 319.8 0.1451 0.4792
1.5 0.7931 382.4 0.1448 0.4834
2.0 0.8063 785.5 0.1409 0.4869
@@ 6 Conclusion
In this paper, we introduce PaSa, a novel paper search agent designed to provide comprehensive and accurate results for complex academic queries. PaSa is implemented within the AGILE, a reinforcement learning framework for LLM agents. To train PaSa, we developed AutoScholarQuery, a dataset of fine-grained academic queries and corresponding papers drawn from top-tier AI conference publications. To evaluate PaSa in real-world scenarios, we also constructed RealScholarQuery, a dataset of actual academic queries paired with annotated papers. Our experimental results demonstrate that PaSa outperforms all baselines, including Google, Google Scholar, and Google with GPT-4o, ChatGPT, GPT-o1, and PaSa-GPT-4o. In particular, PaSa-7B surpasses Google with GPT-4o by 37.78% in recall@20 and 39.90% in recall@50, while also exceeding PaSa-GPT-4o by 30.36% in recall and 4.25% in precision. These findings underscore that PaSa significantly improves the efficiency and accuracy of academic search.
Limitations
(1) Our dataset collection and experiments were primarily focused on the field of machine learning. Although our proposed method is general, we did not explore its performance in other scientific fields. We leave to investigate its applicability to other domains in future work.
(2) Due to resource constraints, our experiments primarily use LLMs with 7b parameters. We expect that scaling up to larger models will lead to more powerful agents. Expanding PaSa to leverage larger LLMs is our future work.
Acknowledgments
The authors thank Yaohua Fang, Zheng Li, Qiang Lu, Yelong Shi, Xuguang Wei, and Tingshuai Yan from ByteDance for their support in developing the PaSa demo. We also thank Jianghui Xie at ByteDance for her assistance with the release of the PaSa demo. Finally, we thank the anonymous reviewers for their valuable suggestions that helped improve this work.
Appendix A Quality Evaluation of AutoScholarQuery
To assess the quality of AutoScholarQuery, we sampled 100 query-paper pairs and evaluated the rationality and relevance of each query and its corresponding paper. The detailed evaluation criteria are as follows:
  • A qualified query should be a complete and understandable sentence. For example, incomplete or fragmented sentences are not acceptable.
  • A query that misrepresents the meaning of the source paper, leading to irrelevant citations, is not qualified. This includes queries that exaggerate the scope or introduce incorrect conditions.
  • A query is ambiguous if there is insufficient context and additional information is needed. For instance, abbreviations with multiple meanings can create ambiguity, leading to the corresponding citations being incomplete answer lists.
  • An answer paper is considered qualified if it aligns with the requirements of the query. The paper should address all or most of the essential factors that make it a suitable response.
Our quality check found that 94.0% of the queries were qualified. Among them, 93.7% of the corresponding answer papers were also qualified. The primary reason for unqualified papers was inaccurate citations in the source paper.
The prompt for search query generation
@@ Appendix A Quality Evaluation of AutoScholarQuery
Appendix B Annotation details
The annotators of RealScholarQuery include professors from the Department of Computer Science at a top-tier university in China. They are paid \$4 per data entry, which consists of a user query and a research paper. Their task is to determine whether the paper satisfies the query.
B.1 Annotation Instructions
For each <user query, paper> pair, carefully assess whether the paper address the user query. Write your decision and provide a brief explanation (1-2 sentences). Specific guidelines are as follows:
  • You may read the entire paper to determine whether it satisfies the user query.
  • Exclude survey papers unless the user query specifically requests them.
  • All conditions in the user query must be met for the paper to be considered qualified.
B.2 Quality control
The annotation process follows the following quality control measures:
  • Stage 1: Annotators work in batches of 20. Authors review 100% of the annotations. Once the consistency rate on the first pass reaches 90%, the process moves to Stage 2.
  • Stage 2: Annotators work in batches of 50. Authors randomly check 40% of the annotations. If the consistency rate is below 90%, the entire batch is re-annotated and re-checked. Once the batch meets the 90% consistency rate on the first pass, the process moves to Stage 3.
  • Stage 3: Annotators work in batches of 100. Authors randomly check 20% of the annotations. If the consistency rate is below 90%, the entire batch is re-annotated and re-checked.
Two authors conducted the quality control.
Appendix C Example in RealScholarQuery
Table 9 presents an example query and corresponding papers from RealScholarQuery.
Appendix D Implementation Details of the Crawler
D.1 Imitation learning data generation
For expanded sessions starting from $S_{q+p}$, we continue by searching for the generated queries using Google. We then sample papers from the search results and obtain the initial state, which includes both the query and a paper. To build the session trajectory, we examine each sub-section of the paper. If the sub-section references at least one paper in the AutoScholarQuery training set corresponding to the query, the sub-section is selected. Otherwise, the sub-section is selected with a 10% probability to enhance trajectory diversity. The selected sections are filled into the template in Table 11, completing the session trajectory. In total, 9,978 expanded session trajectories are constructed.
D.2 PPO training
During PPO training, each device processes 4 user queries in each step, generating a search session for each user query. Then, 6 expansion sessions are created by randomly sampling 6 papers from the search results. This process is repeated with the expanded citation results, yielding 6 additional expanded sessions. In total, 16 session trajectories are generated per step.
Name Value
$\alpha$ (Equation 1) 1.5
$c(\text{[Search]})$ (Equation 1) 0.1
$c(\text{[Expand]})$ (Equation 1) 0.1
$c(\text{[Stop]})$ (Equation 1) 0.0
$\gamma_{0}$ (Equation 4.2) 1.0
$\gamma_{1}$ (Equation 4.2) 0.1
$\beta$ (Equation 4.2) 0.1
$\epsilon$ (Equation 5, Equation 4.2) 0.2
$\eta$ (Equation 8) 10
learning rate 1e-6
epoch per step 2
forward batch size 1
accumulate batch size 16
NVIDIA H100 GPU 16
policy freezing step 50
total step 250
Table 12: The hyperparameters used in PPO training.
Table 12 lists the hyperparameters used during the training process. Figure 3 depicts the RL training curves, which show a steady increase in return with the training steps, eventually converging after 200 steps.
Appendix E Implementation Details of the Selector
The prompt for paper selection
Table 13: Prompt used by PaSa Selector or GPT-4o to evaluate paper relevance.
Google - - - 0.2101 0.2010 0.1788
Google Scholar - - - 0.0801 0.0739 0.0612
Google with GPT-4o - - - 0.2101 0.2010 0.1788
ChatGPT - 0.0507 0.3046 - - -
GPT-o1 - 0.0374 0.2006 - - -
PaSa-GPT-4o 0.7595 0.1817 0.4488 - - -
PaSa-7b 0.7752 0.1881 0.5328 0.6932 0.6543 0.5494
PaSa-7b-ensemble 0.8244 0.1822 0.5568 0.7041 0.6795 0.5535
Table 14: Results on 100-sample subset of AutoScholarQuery test.
We begin by sampling user queries from the AutoScholarQuery training set. For each user query and one of its corresponding papers in the AutoScholarQuery training set, we prompt GPT-4o to generate a decision token and rationale (see Table 13 for the prompt). We reject any data where the decision token is "False", as this contradicts the AutoScholarQuery label. The remaining data are retained as positive <user query, paper> pairs.
@@ Appendix E Implementation Details of the Selector
Appendix F Selector Test Dataset
@@
$c(a_{t})$ Crawler Crawler Precision Recall
Recall Actions
0 0.8239 1296.3 0.1388 0.4852
0.1 0.7931 382.4 0.1448 0.4834
0.2 0.7478 230.1 0.1489 0.4764
Table 15: Performance of the Crawler trained on different action cost $c(a_{t})$ on AutoScholarQuery test set.
Appendix G Additional Experimental Results
G.1 Results on 100-sample subset of AutoScholarQuery
To ensure a fair comparison with the ChatGPT baseline, which is evaluated on only 100 samples from AutoScholarQuery test, we report the performance of all methods on the same subset in Table 14. The results align with those in Table 4, confirming that PaSa-7b consistently outperforms all baselines.
G.2 Action cost
We incorporate action costs to prevent the agent from taking excessive, unproductive actions. Without such costs, the total number of actions would increase significantly without yielding meaningful outcomes.
The key consideration is the reward coefficient $\alpha$ and the action cost $c(a_{t})$ in Equation 1. In Table 8, we fix $c(a_{t})$ and analyze how varying $\alpha$ affects performance.
Additionally, Table 15 shows how different values of $c(a_{t})$ affect the final performance.
Appendix H Prompt Templates
H.1 Prompts used for data synthesis in AutoScholarQuery
Table 16 presents the prompt template used with GPT-4o to automatically generate AutoScholarQuery. For each paper, we extract its Related Work section, input it into GPT-4o, and use the prompt to generate scholarly queries along with their corresponding paper answers.
H.2 Prompts for baselines
Table 17 presents the search query paraphrasing prompt used for the baseline Google with GPT-4o.
Table 18, 19 and 20 show the prompts used for the ChatGPT baseline (search-enabled GPT-4o), the GPT-o1 baseline and PaSa-GPT-4o, respectively.
The prompt for AutoScholarQuery generation
[…]
Table 16: The prompt used with GPT-4o to automatically synthesize AutoScholarQuery.
The prompt for search query paraphrase
@@ H.2 Prompts for baselines
Table 17: The search query paraphrasing prompt used for the Google with GPT-4o baseline.
The prompt for ChatGPT (search-enabled GPT-4o)
@@
Table 18: The prompt for ChatGPT baseline (search-enabled GPT-4o).
The prompt for GPT-o1
{user_query} You should return arxiv papers. You should provide more than 10 paper you searched in JSON format: {"paper_1": {"title": , "authors": , "link": }, "paper_2": {"title": , "authors": , "link": }}. Do not return any other description.
Table 19: The prompt for GPT-o1 baseline.
The prompt for search session of Crawler
The prompt for the expand session of Crawler
You are an elite researcher in the field of AI, currently conducting research on the [topic] specified below. Your task is to determine if the paper specified below likely contains highly relevant citations for the [topic] and, if so, to identify specific sections where these citations might appear. Task Instructions: 1. Relevance Assessment: Decide if the paper is likely to include citations highly relevant to the given [topic]. Output "Yes" or "No" on the first line. 2. Section Selection: If you answered "Yes" in step 1, identify which sections of the paper are likely to contain these relevant citations. From the list of provided sections, select only those you think may contain relevant citations. If no sections seem relevant even if your answer to step 1 was "Yes," leave this empty. Output the selected sections in JSON format on the second line. [topic]: {user_query} [paper title]: {title} [paper abstract]: {abstract} [paper sections]: {sections} Output Format: Output exactly two lines: 1. The first line: Either "Yes" or "No" based on the relevance assessment. 2. The second line: A JSON string with selected sections, e.g., {{"selected_section_1": section_name_1, "selected_section_2": section_name_2}}. If no sections are selected, output {{}}.
The prompt for Selector
@@
Table 20: The prompts for PaSa-GPT-4o.